# Sparse–Smooth Spatially Varying Coefficient Quantile Regression

### v1

@@ Abstract
We develop a convex framework for spatially varying coefficient quantile regression that, for each predictor, separates a location-invariant global effect from a spatial deviation. An adaptive group penalty selects whether a predictor varies over space, while a graph–Laplacian quadratic promotes spatial continuity of the deviations on irregular networks. The formulation is identifiable via degree-weighted centering and scales with sparse linear algebra. We provide two practical solvers—an ADMM algorithm with closed-form proximal maps for the check loss and a smoothed proximal-gradient scheme based on the Moreau envelope—together with implementation guidance (projection for identifiability, stopping diagnostics, and preconditioning). Under mild conditions on the sampling design, covariates, error density, and graph geometry, we establish selection consistency for the deviation groups, mean-squared error bounds that balance Laplacian bias and stochastic variability, and root-$n$ asymptotic normality for the global coefficients with an oracle property. Simulations mimicking air-pollution applications demonstrate accurate recovery of global vs. local effects and competitive predictive performance under heteroskedastic, heavy-tailed noise. We discuss graph construction, spatially blocked cross-validation (to prevent leakage), and options for robust standard errors under spatial dependence.
@@
Quantile regression , spatially varying coefficients , graph Laplacian , group penalty , selection consistency , spatial cross-validation
@@
Global–local decomposition for spatially varying coefficient quantile regression with adaptive $\ell_{2}$-group selection.
Graph-Laplacian smoothing on irregular networks with degree-weighted centering for identifiability.
Two scalable solvers: ADMM with closed-form proximal operators and a Moreau-smoothed proximal-gradient algorithm.
Selection consistency for spatial deviations and root-$n$ asymptotics for global coefficients (oracle property).
Practical guidance on graph construction, spatially blocked cross-validation, and robust standard errors.
@@ 1 Introduction
Spatial data often display nonstationary relations between a response and its predictors, so that regression coefficients vary over space. The geographically weighted regression (GWR) literature formalizes this idea via locally weighted fits across locations [Fotheringham2002]. Extensions include multiscale GWR to allow distinct spatial scales across coefficients [FotheringhamYangKang2017, OshanFotheringham2019]. In parallel, quantile regression (QR) [KoenkerBassett1978, Koenker2005] models conditional quantiles and is known to be robust under heavy tails and heteroskedasticity. Combining spatially varying coefficients (SVCs) with QR yields a tool for tail behavior and distributional heterogeneity. Bayesian and frequentist SVC–QR formulations have been studied using Gaussian-process priors or nonparametric components [Gelfand2003SVC, Reich2012STQR], but scalable convex formulations that simultaneously deliver interpretable global-vs-local decisions and graph-based smoothness remain relatively limited.
We propose a penalized SVC–QR estimator with two desiderata. First, for each covariate we separate a global component (constant in space) from a spatial deviation, and we use a group $L_{2}$ penalty to shrink the deviation to zero whenever the covariate is effectively global. This provides an interpretable global-vs-local decision not available in kernel-weighted local QR. Second, conditioned on being local, we regularize the deviation by a graph Laplacian quadratic constructed from the sampling locations. This encourages spatial continuity and can accommodate relatively sharp local changes when supported by the data; strictly edge-preserving behavior would require total-variation-type penalties, which we discuss as an extension. The formulation is convex and compatible with large-scale QR solvers, including proximal-gradient implementations under a smooth surrogate and ADMM schemes tailored to the check loss.
@@
On theory, we show that global coefficients are root-$n$ consistent and asymptotically normal, deviation fields converge in mean square at a rate balancing Laplacian bias and stochastic variability, and selection of globally constant effects is consistent under a groupwise beta-min condition. Our assumptions allow random sampling locations filling a compact set, bounded covariates, and sparse graphs with uniformly bounded average degree or controlled growth.
We emphasize choices for reproducibility: mutual $k$-nearest-neighbor or compact-kernel graphs, spatially blocked cross-validation (to mitigate leakage), and diagnostics for KKT conditions and stopping. We also document computational scaling and preconditioning for large geospatial datasets.
@@ 2 Model
We develop a sparse–smooth spatially varying coefficient quantile regression (SS–SVCQR) at a fixed quantile $\tau\in(0,1)$. Scalars are in italic lower case, vectors in bold lower case, and matrices in bold upper case. For $\bm{v}$, $\|\bm{v}\|_{2}$ is the Euclidean norm, and $\bm{1}$ is an all-ones vector of conformable length.
We observe $\{(Y_{i},\bm{Z}_{i},\bm{X}_{i},u_{i}):i=1,\dots,n\}$, with $Y_{i}\in\mathbb{R}$, $\bm{Z}_{i}\in\mathbb{R}^{q}$ (purely global effects), $\bm{X}_{i}\in\mathbb{R}^{p}$ (potentially varying over space), and $u_{i}\in\mathcal{U}\subset\mathbb{R}^{2}$ the location. Let $Q_{\tau}(Y_{i}\mid\bm{Z}_{i},\bm{X}_{i},u_{i})$ denote the conditional $\tau$-quantile. The target relation is
$Q_{\tau}\{Y_{i}\mid\bm{Z}_{i},\bm{X}_{i},u_{i}\}=\bm{Z}_{i}^{\top}\bm{\alpha}+\sum_{j=1}^{p}X_{ij}\beta_{j}(u_{i}),\qquad i=1,\dots,n,$ (1)
with $\bm{\alpha}\in\mathbb{R}^{q}$ and spatially varying coefficients $\beta_{j}(\cdot)$. Each coefficient is decomposed into a global level and a spatial deviation,
$\beta_{j}(u)=\beta_{G,j}+\delta_{j}(u),\qquad j=1,\dots,p,$ (2)
where $\bm{\beta}_{G}=(\beta_{G,1},\dots,\beta_{G,p})^{\top}\in\mathbb{R}^{p}$ and $\delta_{j}:\mathcal{U}\to\mathbb{R}$.
@@
$\Pr\big\{Y_{i}\leq\bm{Z}_{i}^{\top}\bm{\alpha}_{0}+\sum_{j=1}^{p}X_{ij}\beta_{0,j}(u_{i})\ \big|\ \bm{Z}_{i},\bm{X}_{i},u_{i}\big\}=\tau\quad\text{almost surely for each }i.$ (3)
@@
Spatial geometry is encoded by a symmetric graph on $\{u_{i}\}$. Let $\bm{A}\in\mathbb{R}^{n\times n}$ be a symmetric adjacency from a mutual $k$-nearest-neighbor rule or a symmetric compact-support kernel, and let $\bm{D}=\operatorname{diag}(\bm{A}\bm{1})$ collect node degrees. We use the symmetric normalized Laplacian
$\bm{L}_{\mathrm{sym}}=\bm{I}-\bm{D}^{-1/2}\bm{A}\bm{D}^{-1/2},$ (4)
which is symmetric positive semidefinite with spectrum in $[0,2]$.
To make (2) unique and prevent drift between global and local parts, local deviations are degree-weight centered componentwise. For a component $\mathcal{C}\subset\{1,\dots,n\}$, let $\bm{1}_{\mathcal{C}}$ be its indicator. Writing $\bm{\delta}_{j}=(\delta_{j}(u_{1}),\dots,\delta_{j}(u_{n}))^{\top}$, the identifiability constraint is
$\bm{1}_{\mathcal{C}}^{\top}\bm{D}\bm{\delta}_{j}=0$ (5)
for every connected component $\mathcal{C}$ and all $j$. Because $\mathrm{null}(\bm{L}_{\mathrm{sym}})$ on a component is spanned by $\bm{D}^{1/2}\bm{1}_{\mathcal{C}}$, orthogonality of $\bm{D}^{1/2}\bm{\delta}_{j}$ to the null space is equivalent to (5). Under (5), $\beta_{G,j}$ equals the degree-weighted average of the spatial coefficient $\beta_{j}(\cdot)$ across nodes; $\delta_{j}(\cdot)$ captures pure spatial departures around that average.
Vectorization at observed sites fixes the linear operators. Stack $\bm{y}=(Y_{1},\dots,Y_{n})^{\top}$, $\bm{Z}=[\bm{Z}_{1}^{\top};\cdots;\bm{Z}_{n}^{\top}]\in\mathbb{R}^{n\times q}$, and $\bm{X}=[\bm{X}_{1}^{\top};\dots;\bm{X}_{n}^{\top}]\in\mathbb{R}^{n\times p}$. For $j=1,\dots,p$, define $\bm{X}_{\odot j}=\operatorname{diag}(X_{1j},\dots,X_{nj})\in\mathbb{R}^{n\times n}$ and $\bm{\delta}_{j}\in\mathbb{R}^{n}$. Concatenate deviations by
$\bm{\Delta}=\big[\bm{\delta}_{1}^{\top},\ \bm{\delta}_{2}^{\top},\ \dots,\ \bm{\delta}_{p}^{\top}\big]^{\top}\in\mathbb{R}^{np},$
and build the block operator
@@
The conditional quantile vector implied by (1)–(2) is
$\bm{q}_{\tau}=\bm{Z}\bm{\alpha}+\bm{X}\bm{\beta}_{G}+\bm{X}_{\delta}\bm{\Delta},$ (7)
and the residuals at $\theta=(\bm{\alpha},\bm{\beta}_{G},\bm{\Delta})$ are $\bm{r}(\theta)=\bm{y}-\bm{q}_{\tau}$. The local component $\bm{X}_{\delta}\bm{\Delta}=\sum_{j=1}^{p}\bm{X}_{\odot j}\bm{\delta}_{j}$ is a superposition of $p$ diagonal-scaled fields.
Spatial smoothness is summarized by
$\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j}=\tfrac{1}{2}\sum_{(i,\ell)\in\mathcal{E}}A_{i\ell}\Big\{\tfrac{\delta_{j}(u_{i})}{\sqrt{d_{i}}}-\tfrac{\delta_{j}(u_{\ell})}{\sqrt{d_{\ell}}}\Big\}^{2},\qquad d_{i}=(\bm{D})_{ii},$ (8)
where $\mathcal{E}$ is the set of undirected edges (each unordered pair counted once). The quantity vanishes only for degree-weighted constants on each component.
Interpretation follows: a predictor $X_{:j}$ is purely global when $\bm{\delta}_{j}=\bm{0}$ and $\beta_{G,j}\neq 0$, purely local with zero degree-weighted mean when $\beta_{G,j}=0$ and $\bm{\delta}_{j}\neq\bm{0}$, and mixed when both parts are nonzero. Disconnected graphs are handled componentwise without changing notation.
@@ 3 Estimation and Algorithms
$\min_{\bm{\theta}_{P},\bm{\Theta}_{L}}\sum_{i=1}^{n}\rho_{\tau}\Big(Y_{i}-\bm{Z}_{i}^{\top}\bm{\alpha}-\bm{X}_{i}^{\top}\bm{\beta}_{G}-\sum_{j=1}^{p}X_{ij}\delta_{j}(u_{i})\Big)+\lambda_{1}\sum_{j=1}^{p}w_{j}\|\bm{\delta}_{j}\|_{2}+\lambda_{2}\sum_{j=1}^{p}\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j},$ (9)
subject to $\bm{1}_{\mathcal{C}}^{\top}\bm{D}\bm{\delta}_{j}=0$ for every component $\mathcal{C}$ and all $j$. Weights $w_{j}>0$ enable adaptive grouping, e.g. $w_{j}=(\|\tilde{\bm{\delta}}_{j}\|_{2}+a)^{-\gamma}$ with $a>0$ and $\gamma\in(0,1]$. Tuning $(\lambda_{1},\lambda_{2})$ is selected by spatially blocked cross-validation.
@@
subject to $\bm{s}=\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{j=1}^{p}\bm{X}_{\odot j}\bm{\delta}_{j}$ and $\bm{z}_{j}=\bm{\delta}_{j}$, together with centering via (10). Using scaled duals $(\bm{u},\{\bm{v}_{j}\})$ and penalties $(\rho_{s},\rho_{z})>0$, one ADMM sweep is:
@@
$\left[\begin{array}[]{cc}\bm{Z}^{\top}\bm{Z}&\bm{Z}^{\top}\bm{X}\\ \bm{X}^{\top}\bm{Z}&\bm{X}^{\top}\bm{X}\end{array}\right]\left[\begin{array}[]{c}\bm{\alpha}\\ \bm{\beta}_{G}\end{array}\right]=\left[\begin{array}[]{c}\bm{Z}^{\top}(\bm{y}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}-\bm{s}+\bm{u})\\[2.0pt] \bm{X}^{\top}(\bm{y}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}-\bm{s}+\bm{u})\end{array}\right].$
@@
$\mathrm{prox}_{\gamma\rho_{\tau}}(v)=\begin{cases}v-\gamma\tau,&v>\gamma\tau,\\ 0,&-\gamma(1-\tau)\leq v\leq\gamma\tau,\\ v+\gamma(1-\tau),&v<-\gamma(1-\tau).\end{cases}$
@@
$\big(2\lambda_{2}\bm{L}_{\mathrm{sym}}+\rho_{s}\bm{X}_{\odot j}^{\top}\bm{X}_{\odot j}+\rho_{z}\bm{I}\big)\bm{\delta}_{j}=\rho_{s}\bm{X}_{\odot j}^{\top}\big(\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{\ell\neq j}\bm{X}_{\odot\ell}\bm{\delta}_{\ell}-\bm{s}+\bm{u}\big)+\rho_{z}(\bm{z}_{j}-\bm{v}_{j}).$ (13)
@@
$\bm{z}_{j}\leftarrow\Big(1-\frac{\lambda_{1}w_{j}}{\rho_{z}\|\bm{\delta}_{j}+\bm{v}_{j}\|_{2}}\Big)_{+}(\bm{\delta}_{j}+\bm{v}_{j}),\qquad j=1,\dots,p.$ (14)
@@
$\bm{u}\leftarrow\bm{u}+\big(\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}-\bm{s}\big),\qquad\bm{v}_{j}\leftarrow\bm{v}_{j}+(\bm{\delta}_{j}-\bm{z}_{j}),$
@@
$\bm{r}_{s}=\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}-\bm{s},\qquad\bm{r}_{z,j}=\bm{\delta}_{j}-\bm{z}_{j}$
@@
$\bm{d}_{s}=\rho_{s}(\bm{s}-\bm{s}^{\mathrm{prev}}),\qquad\bm{d}_{z,j}=\rho_{z}(\bm{z}_{j}-\bm{z}_{j}^{\mathrm{prev}})$

### v2

@@ Abstract
Modeling spatially varying relationships in the presence of heavy-tailed responses and irregular sampling designs presents a significant challenge in geospatial analysis. In this paper, we propose a Sparse–Smooth Spatially Varying Coefficient Quantile Regression (SS–SVCQR) framework that simultaneously achieves robust estimation, structural interpretability, and computational scalability. Our approach decomposes each coefficient function into a global baseline and a spatial deviation component. By applying a group $L_{2}$ penalty, the model automatically selects between globally constant and spatially varying specifications, offering a parsimonious alternative to standard Geographically Weighted Regression (GWR). Conditional on local variation, spatial smoothness is enforced via a normalized graph Laplacian regularization, which adapts to varying sampling densities and admits a Bayesian interpretation as an intrinsic Gaussian Markov Random Field (GMRF) prior. We develop efficient convex optimization algorithms based on ADMM and smoothed proximal-gradient methods to handle the non-differentiable check loss. Theoretically, we establish the root-$n$ consistency and asymptotic normality of the global estimators, as well as the selection consistency of the global-vs-local structure under mild conditions. Extensive simulations and a real-world housing application show that SS–SVCQR provides more accurate quantile predictions and more interpretable global–local structures than existing kernel- and spline-based alternatives, especially in the presence of heavy-tailed noise.
@@
Quantile Regression , Spatially Varying Coefficients , Graph Laplacian , Variable Selection , Group Lasso , ADMM.
@@
Parsimonious global–local decomposition distinguishes constant effects from spatial deviations via adaptive group selection.
Normalized Graph-Laplacian smoothing adapts to irregular sampling densities while ensuring geometric fidelity.
Convex formulation solved by efficient ADMM and Moreau-smoothed proximal-gradient algorithms.
Establishes oracle properties: selection consistency for local deviations and root-$n$ normality for global coefficients.
Outperforms kernel-based methods in recovering heterogeneous patterns under heavy-tailed distributions.
@@ 1 Introduction
Spatial nonstationarity—the phenomenon where relationships between a response variable and its predictors vary across geographical locations—poses a fundamental challenge in regression analysis. Ignoring this heterogeneity by fitting a global constant-coefficient model can lead to biased estimates and misleading inference. To address this, the literature on Spatially Varying Coefficient (SVC) models has flourished. The most prominent approach is Geographically Weighted Regression (GWR) [Brunsdon1996, Fotheringham2002], which fits local weighted regressions at each target location. While effective, standard GWR often suffers from multicollinearity and bandwidth selection sensitivity. Recent extensions have sought to mitigate these issues through multiscale formulations [FotheringhamYangKang2017, OshanFotheringham2019] and mixed GWR (MGWR) that allows some coefficients to remain fixed [Mei2006, Geniaux2017]. Alternative SVC frameworks include Bayesian hierarchical models based on Gaussian Processes (GPs) [Gelfand2003SVC, Banerjee2008] and frequentist approaches using penalized splines [Eilers2010, Xiao2013]. However, GP-based methods often face cubic computational complexity, limiting their applicability to large-scale geospatial datasets.
Parallel to mean regression, Quantile Regression (QR) [KoenkerBassett1978] has become an indispensable tool for characterizing the entire conditional distribution. In spatial contexts, modeling conditional quantiles is particularly valuable because spatial heterogeneity often manifests asymmetrically—affecting the tails of the distribution differently than the center [Reich2012STQR]. Although Spatially Varying Coefficient Quantile Regression (SVC-QR) offers a powerful lens for such analysis, existing methods face dual challenges: interpretability and scalability. Many nonparametric SVC-QR estimators [Hallin2004, Lum2012] yield coefficients that vary everywhere, failing to distinguish between truly local effects and those that are effectively global. This lack of parsimony complicates interpretation. While variable selection techniques like the Lasso have been adapted for spatial models [Wang2008, Cai2013], scalable convex formulations that simultaneously deliver interpretable global-vs-local selection and graph-based smoothness remain relatively limited. Recent developments in ”Network Lasso” and graph-trend filtering [Hallac2015, Wang2016] suggest that graph-based penalties offer a computationally efficient alternative to kernel or spline methods, yet their integration with quantile loss in a spatial setting, together with an explicit mechanism for global–vs–local selection, remains underexplored.
In this paper, we propose a penalized estimator that addresses these gaps by integrating graph Laplacian regularization with group-sparse penalization. Our approach, termed SS-SVCQR (Sparse-Smooth Spatially Varying Coefficient Quantile Regression), is designed with two key desiderata. First, we employ a decomposition strategy that separates each coefficient into a global baseline and a spatial deviation. A group $L_{2}$ penalty [YuanLin2006] is applied to the deviation vector, shrinking it to zero when the covariate effect is globally constant. This provides an automated, interpretable decision mechanism for global-vs-local specification, a feature often absent in kernel-weighted local QR. Second, conditioned on being local, we regularize the deviations using a normalized graph Laplacian constructed from the sampling locations [Smola2003, Belkin2003]. This encourages spatial continuity and allows the model to adapt to irregular sampling designs without the heavy computational burden of dense covariance matrices used in GPs. The resulting formulation is convex and compatible with large-scale solvers, including Alternating Direction Method of Multipliers (ADMM) schemes tailored to the non-differentiable check loss.
@@
On the theoretical front, we establish that our estimator enjoys desirable asymptotic properties. We show that the global coefficients are root-$n$ consistent and asymptotically normal, while the deviation fields converge in mean square at a rate that balances Laplacian bias and stochastic variability. Furthermore, we prove that the selection of globally constant effects is consistent under a groupwise beta-min condition. Our theoretical framework accommodates random sampling locations filling a compact set and sparse graphs with uniformly bounded average degrees.
The remainder of this paper is organized as follows. Section 2 introduces the proposed model decomposition and the graph Laplacian regularization framework. Section 3 details the optimization algorithms, including an efficient ADMM implementation and tuning parameter selection via spatially blocked cross-validation. Section 4 provides the asymptotic analysis. In Section 5, we evaluate the finite-sample performance of our method against competing GWR and spline-based approaches through extensive simulations. Section 6 applies our method to a real-world analysis of housing prices in the Lucas County, demonstrating the model’s ability to uncover heterogeneous pricing mechanisms and distributional tail dynamics. Section 7 concludes with a discussion.
@@ 2 Model
We develop a sparse–smooth spatially varying coefficient quantile regression (SS–SVCQR) at a fixed quantile $\tau\in(0,1)$. This framework fuses the robustness of quantile regression with the flexibility of graph-based regularization to capture heterogeneous spatial dependencies.
Scalars are denoted by italic lower case, vectors by bold lower case, and matrices by bold upper case. For a vector $\bm{v}$, $\|\bm{v}\|_{2}$ denotes the Euclidean norm, and $\bm{1}$ is an all-ones vector of conformable length.
2.1 Data and Decomposition Structure
We observe a spatial dataset $\{(Y_{i},\bm{Z}_{i},\bm{X}_{i},u_{i}):i=1,\dots,n\}$, where $Y_{i}\in\mathbb{R}$ is the response, $\bm{Z}_{i}\in\mathbb{R}^{q}$ represents covariates with purely global effects, $\bm{X}_{i}\in\mathbb{R}^{p}$ contains predictors with potentially spatially varying effects, and $u_{i}\in\mathcal{U}\subset\mathbb{R}^{2}$ denotes the spatial location. Let $Q_{\tau}(Y_{i}\mid\bm{Z}_{i},\bm{X}_{i},u_{i})$ be the conditional $\tau$-quantile. We assume the data generating process follows a varying-coefficient structure:
$Q_{\tau}\{Y_{i}\mid\bm{Z}_{i},\bm{X}_{i},u_{i}\}=\bm{Z}_{i}^{\top}\bm{\alpha}+\sum_{j=1}^{p}X_{ij}\beta_{j}(u_{i}),\quad i=1,\dots,n,$ (1)
where $\bm{\alpha}\in\mathbb{R}^{q}$ collects the fixed global coefficients, and $\beta_{j}(\cdot)$ are unknown functions characterizing the spatial heterogeneity. To simultaneously perform variable selection between global and local specifications, we decompose each varying coefficient into a global baseline level and a location-specific spatial deviation:
$\beta_{j}(u)=\beta_{G,j}+\delta_{j}(u),\quad j=1,\dots,p.$ (2)
Here, $\bm{\beta}_{G}=(\beta_{G,1},\dots,\beta_{G,p})^{\top}\in\mathbb{R}^{p}$ captures the domain-wide average effect, while $\delta_{j}:\mathcal{U}\to\mathbb{R}$ captures local departures. This additive form allows the model to adaptively shrink $\delta_{j}(\cdot)$ to zero when the effect of predictor $j$ is homogeneous, reducing the model to a standard global quantile regression.
@@ 2.1 Data and Decomposition Structure
$\Pr\big\{Y_{i}\leq\bm{Z}_{i}^{\top}\bm{\alpha}_{0}+\sum_{j=1}^{p}X_{ij}\beta_{0,j}(u_{i})\big|\bm{Z}_{i},\bm{X}_{i},u_{i}\big\}=\tau,\ \text{almost surely for each }i.$ (3)
@@
2.2 Graph Construction and Spatial Regularization
Since the locations $\{u_{i}\}$ may be irregularly distributed, we encode the spatial domain’s geometry via a proximity graph. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with vertices $\mathcal{V}=\{1,\dots,n\}$ representing the observations. We construct a sparse, symmetric adjacency matrix $\bm{A}\in\mathbb{R}^{n\times n}$ using a mutual $k$-nearest-neighbor ($k$-NN) rule combined with a Gaussian kernel. Specifically, for two locations $u_{i}$ and $u_{\ell}$, the edge weight is defined as:
$A_{i\ell}=\begin{cases}\exp\left(-\frac{\|u_{i}-u_{\ell}\|_{2}^{2}}{\sigma^{2}}\right),&\text{if }u_{\ell}\in\mathcal{N}_{k}(u_{i})\text{ or }u_{i}\in\mathcal{N}_{k}(u_{\ell}),\\ 0,&\text{otherwise},\end{cases}$ (4)
where $\mathcal{N}_{k}(u)$ denotes the set of $k$ nearest neighbors of $u$, and $\sigma$ is a bandwidth parameter. This construction ensures the graph is connected while adapting to local sampling densities.
To enforce spatial smoothness on $\bm{\delta}_{j}$, we utilize the spectral properties of the graph. Let $\bm{D}=\operatorname{diag}(\bm{A}\bm{1})$ be the degree matrix. We adopt the symmetric normalized Laplacian:
$\bm{L}_{\mathrm{sym}}=\bm{I}-\bm{D}^{-1/2}\bm{A}\bm{D}^{-1/2}.$ (5)
We favor $\bm{L}_{\mathrm{sym}}$ over the unnormalized Laplacian ($\bm{D}-\bm{A}$) because its spectrum is bounded in $[0,2]$ and, crucially, it normalizes for node degree. In spatial statistics, sampling locations often exhibit clustering; the normalized Laplacian ensures that the regularization penalty is not artificially inflated in densely sampled regions [Chung1997].
The spatial roughness of the deviation vector $\bm{\delta}_{j}=(\delta_{j}(u_{1}),\dots,\delta_{j}(u_{n}))^{\top}$ is quantified by the quadratic form:
$\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j}=\frac{1}{2}\sum_{(i,\ell)\in\mathcal{E}}A_{i\ell}\left\{\frac{\delta_{j}(u_{i})}{\sqrt{d_{i}}}-\frac{\delta_{j}(u_{\ell})}{\sqrt{d_{\ell}}}\right\}^{2},\quad d_{i}=(\bm{D})_{ii}.$ (6)
This penalty term admits a probabilistic interpretation. It corresponds to placing an intrinsic Gaussian Markov Random Field (GMRF) prior on the spatial deviations, $p(\bm{\delta}_{j})\propto\exp(-\frac{\lambda_{2}}{2}\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j})$. This prior encodes the belief that, conditional on its neighbors, the deviation at a node is normally distributed around the degree-weighted average of its neighbors, thereby enforcing local coherence.
2.3 Identifiability and Matrix Formulation
The decomposition in (2) requires a constraint to prevent collinearity between the global intercept $\beta_{G,j}$ and the spatial field $\delta_{j}(\cdot)$. We impose that local deviations are degree-weight centered componentwise. For a connected component $\mathcal{C}\subset\{1,\dots,n\}$ with an indicator vector $\bm{1}_{\mathcal{C}}$, the identifiability constraint is:
$\bm{1}_{\mathcal{C}}^{\top}\bm{D}\bm{\delta}_{j}=0,$ (7)
for all $j=1,\dots,p$. Since the null space of $\bm{L}_{\mathrm{sym}}$ on a component is spanned by $\bm{D}^{1/2}\bm{1}_{\mathcal{C}}$, ensuring $\bm{D}^{1/2}\bm{\delta}_{j}$ is orthogonal to this null space is equivalent to (7). Under this constraint, $\beta_{G,j}$ interprets as the degree-weighted spatial average of the total coefficient $\beta_{j}(\cdot)$, while $\delta_{j}(\cdot)$ captures pure spatial variation.
For estimation, we vectorize the system. Stack the responses $\bm{y}=(Y_{1},\dots,Y_{n})^{\top}$ and design matrices $\bm{Z}=[\bm{Z}_{1}^{\top};\cdots;\bm{Z}_{n}^{\top}]\in\mathbb{R}^{n\times q}$ and $\bm{X}=[\bm{X}_{1}^{\top};\dots;\bm{X}_{n}^{\top}]\in\mathbb{R}^{n\times p}$. To handle the varying coefficients, we define diagonal operators $\bm{X}_{\odot j}=\operatorname{diag}(X_{1j},\dots,X_{nj})\in\mathbb{R}^{n\times n}$ for $j=1,\dots,p$. Concatenating the deviations as $\bm{\Delta}=\big[\bm{\delta}_{1}^{\top},\dots,\bm{\delta}_{p}^{\top}\big]^{\top}\in\mathbb{R}^{np}$, we construct the block operator:
@@ 2.3 Identifiability and Matrix Formulation
so that $X_{\delta}\Delta$ stacks all local deviation terms in a linear operator form. The global vector of conditional quantiles is then succinctly expressed as:
$\bm{q}_{\tau}=\bm{Z}\bm{\alpha}+\bm{X}\bm{\beta}_{G}+\bm{X}_{\delta}\bm{\Delta}.$ (9)
This linear formulation allows us to leverage efficient convex optimization algorithms to estimate the parameters $\theta=(\bm{\alpha},\bm{\beta}_{G},\bm{\Delta})$ by minimizing the check loss subject to the penalties defined above.
@@ 3 Estimation and Algorithms
$\min_{\bm{\theta}_{P},\bm{\Theta}_{L}}\sum_{i=1}^{n}\rho_{\tau}\Big(Y_{i}-\bm{Z}_{i}^{\top}\bm{\alpha}-\bm{X}_{i}^{\top}\bm{\beta}_{G}-\sum_{j=1}^{p}X_{ij}\delta_{j}(u_{i})\Big)+\lambda_{1}\sum_{j=1}^{p}w_{j}\|\bm{\delta}_{j}\|_{2}$ (10)
$+\lambda_{2}\sum_{j=1}^{p}\bm{\delta}_{j}^{\top}\bm{L}_{\mathrm{sym}}\bm{\delta}_{j},$
subject to $\bm{1}_{\mathcal{C}}^{\top}\bm{D}\bm{\delta}_{j}=0$ for every component $\mathcal{C}$ and all $j$. Weights $w_{j}>0$ enable adaptive grouping, e.g. $w_{j}=(\|\tilde{\bm{\delta}}_{j}\|_{2}+a)^{-\gamma}$ with $a>0$ and $\gamma\in(0,1]$[Zou2006, WangLeng2008]. Tuning $(\lambda_{1},\lambda_{2})$ is selected by spatially blocked cross-validation.
@@
subject to $\bm{s}=\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{j=1}^{p}\bm{X}_{\odot j}\bm{\delta}_{j}$ and $\bm{z}_{j}=\bm{\delta}_{j}$, together with centering via (11). Using scaled duals $(\bm{u},\{\bm{v}_{j}\})$ and penalties $(\rho_{s},\rho_{z})>0$, an ADMM sweep is as follows[Boyd2011]:
@@
$\begin{bmatrix}\bm{Z}^{\top}\bm{Z}&\bm{Z}^{\top}\bm{X}\\ \bm{X}^{\top}\bm{Z}&\bm{X}^{\top}\bm{X}\\ \end{bmatrix}\begin{bmatrix}\bm{\alpha}\\ \bm{\beta}_{G}\\ \end{bmatrix}=\begin{bmatrix}\bm{Z}^{\top}(\bm{y}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}-\bm{s}+\bm{u})\\ \bm{X}^{\top}(\bm{y}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}-\bm{s}+\bm{u})\\ \end{bmatrix}$.
@@
[…]
@@
$\big(2\lambda_{2}\bm{L}_{\mathrm{sym}}+\rho_{s}\bm{X}_{\odot j}^{\top}\bm{X}_{\odot j}+\rho_{z}\bm{I}\big)\bm{\delta}_{j}=\rho_{s}\bm{X}_{\odot j}^{\top}\big(\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{\ell\neq j}\bm{X}_{\odot\ell}\bm{\delta}_{\ell}-\bm{s}+\bm{u}\big)$ (14)
$+\rho_{z}(\bm{z}_{j}-\bm{v}_{j}).$
@@
$\bm{z}_{j}\leftarrow\Big(1-\frac{\lambda_{1}w_{j}}{\rho_{z}\|\bm{\delta}_{j}+\bm{v}_{j}\|_{2}}\Big)_{+}(\bm{\delta}_{j}+\bm{v}_{j}),\quad j=1,\dots,p.$ (15)
@@
$\bm{u}$ $\leftarrow\bm{u}+\big(\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}-\bm{s}\big),$
$\bm{v}_{j}$ $\leftarrow\bm{v}_{j}+(\bm{\delta}_{j}-\bm{z}_{j}),$
@@
$\bm{r}_{s}=\bm{y}-\bm{Z}\bm{\alpha}-\bm{X}\bm{\beta}_{G}-\sum_{j}\bm{X}_{\odot j}\bm{\delta}_{j}-\bm{s},\quad\bm{r}_{z,j}=\bm{\delta}_{j}-\bm{z}_{j}$
@@
$\bm{d}_{s}=\rho_{s}(\bm{s}-\bm{s}^{\mathrm{prev}}),\quad\bm{d}_{z,j}=\rho_{z}(\bm{z}_{j}-\bm{z}_{j}^{\mathrm{prev}})$